# Context-Aware Code Wiring Recommendation with LLM-based Agent

### v1

@@ Abstract
Copy-paste-modify is a widespread and pragmatic practice in software development, where developers adapt reused code snippets—sourced from platforms such as Stack Overflow, GitHub, or LLM outputs—into their local codebase. A critical yet underexplored aspect of this adaptation is code wiring, which involves substituting unresolved variables in the pasted code with suitable ones from the surrounding context. Existing solutions either rely on heuristic rules or historical templates, often failing to effectively utilize contextual information, despite studies showing that over half of adaptation cases are context-dependent. In this paper, we introduce WIRL, an LLM-based agent for code wiring framed as a Retrieval-Augmented Generation (RAG) infilling task. WIRL combines an LLM, a customized toolkit, and an orchestration module to identify unresolved variables, retrieve context, and perform context-aware substitutions. To balance efficiency and autonomy, the agent adopts a mixed strategy: deterministic rule-based steps for common patterns, and a state-machine-guided decision process for intelligent exploration. We evaluate WIRL on a carefully curated, high-quality dataset consisting of real-world code adaptation scenarios. Our approach achieves an exact match precision of 91.7% and a recall of 90.0%, outperforming advanced LLMs by 22.6 and 13.7 percentage points in precision and recall, respectively, and surpassing IntelliJ IDEA by 54.3 and 49.9 percentage points. These results underscore its practical utility, particularly in contexts with complex variable dependencies or multiple unresolved variables. We believe WIRL paves the way for more intelligent and context-aware developer assistance in modern IDEs.
Index Terms:
Copy-paste-modify Practices, Code Reuse, Code Adaptation, Large Language Models, Agent.
@@ I Introduction
Copy-paste-modify is a common and inevitable practice during development. The developers frequently reuse code snippets from online programming Q&A communities, e.g., Stack Overflow [16], open-source repositories, e.g., GitHub [15], or code generation of LLMs. However, in most cases the reused code snippets need additional adaptations to integrate to the local codebase. As reported, more than 85% code snippets from Stack Overflow need to be adapted (modified) before integration into the local code [46, 47]. Variable identifiers are usually under adaptation since undeclared or conflict identifiers would lead to compilation errors or potential bugs. Consequently, code wiring, replacing unresolved variables with existing ones from the local context, is one of the most prevalent forms of adaptation. This term aligns with the analogy of electrical wiring, where components (variables in code) are intentionally connected to establish functionality. An example of code wiring from Stack overflow [9] to GitHub [10] is presented in Fig. 1. Automatic code wiring can help developers get rid of the repetitive and trivial processes during the adaptation and concentrate on the complex business logic. However, existing approaches are designed based on either simple heuristic rules [6] or templates extracted from historical modifications [12], leaving the context not effectively leveraged although 56.1% of the adaptation cases are dependent on the surrounding context as reported by Zhang et al. [46]. Full-parameter LLMs have demonstrated promising performance; however, their high latency renders them impractical for real-world applications. On the other hand, distilled LLMs with smaller parameter sizes tend to under perform in terms of accuracy and robustness.
@@
To this end, we present WIRL, an LLM-based agent for code Wiring through Infilling with RAG and LLMs. WIRL consists of three core components: an LLM, a customized toolkit, and an agent pilot. The customized toolkit offers three primary functionalities: (1) identifying and locating unresolved elements, (2) analyzing and collecting contextual information, and (3) infilling and recommending suitable substitutions. The agent pilot is responsible not only for initializing the prompt but also for coordinating the interaction between the LLM agent and the toolkit. It parses the LLM’s output, updates the prompt dynamically, and invokes the appropriate tools as needed. With a dynamically updated prompt, WIRL incrementally gathers relevant context by invoking appropriate tools until the agent determines that sufficient information has been collected to make a final recommendation. To fully leverage the capabilities of LLMs, we reformulate the adaptation task as a retrieval-augmented generation (RAG)-based infilling task for the unresolved elements—framing it as a code completion problem, which aligns more naturally with the strengths of LLMs. Empirical evidence from Zhang et al. [30] supports this reformulation, showing that LLM performance on code snippet adaptation tasks lags behind their performance on code completion and generation tasks. Moreover, by adopting the RAG-based infilling strategy, even distilled LLMs with smaller parameter sizes can outperform full-parameter LLMs, thereby meeting both the latency and accuracy requirements of real-world software development scenarios. To enhance the efficiency of WIRL, we adopt a hybrid execution mode for the agent, wherein essential steps are extracted and executed initially without invoking the LLM. In addition to this optimization, we introduce a state machine to guide the agent’s exploration process, thereby reducing unnecessary or ineffective attempts and improving overall execution efficiency.
@@
  • We develop a customized toolkit that provides essential functionalities, enabling the agent to perform precise and efficient code adaptation.
@@
The rest of this paper is structured as follows. Section II motivates this study. Section III illustrates the design of WIRL. Section IV presents experimental details and results. Section V discusses the threats and limitations of our study. Section VI discusses the related works, and Section VII concludes this paper.
@@ II Motivating Example
Before illustrating how WIRL operates, we first present the problem formulation, which serves as an important foundation of our approach. Assume the set of unresolved elements is denoted by $\mathcal{U}={u_{i}}$ and the set of context elements by $\mathcal{C}={c_{i}}$. The code wiring task can be formally defined as identifying an injective mapping $\mathcal{M}:\mathcal{U}\rightarrow\mathcal{C}$, where each $u_{i}\in\mathcal{U}$ is mapped to a distinct $c_{i}\in\mathcal{C}$. In this work, however, we reformulate code wiring as a retrieval-augmented generation (RAG)-based infilling task. Specifically, each $u_{i}\in\mathcal{U}$ is replaced with a placeholder token “<Infill>” in the isolated code. The objective then becomes to infill these placeholders using the retrieved context information, including the variable names themselves, which is motivated by the observation that literal similarity often provides valuable contextual cues. This reformulation is grounded in the insight that code completion is more naturally aligned with the capabilities of large language models (LLMs) than explicit code adaptation. This design choice is further supported by recent empirical findings by Zhang et al. [30], which highlight the superior performance of LLMs in code completion tasks relative to direct adaptation scenarios.
@@
Figure 1: Motivating Example
@@ III-A Overview
Figure 2: Overview of WIRL
The overview of WIRL is illustrated in Fig. 2. WIRL consists of three core components: an LLM, a customized toolkit containing three types of tools, i.e., locator, collector, and completer, and an agent pilot, which orchestrates communication between the LLM and the toolkit by parsing outputs, updating prompts, and executing tool invocations. Given a piece of isolated code, the agent pilot begins by initializing the prompt (Step 1). This prompt includes system roles, task descriptions, and instructions on how the task should be addressed using the available tools. It also embeds the location of unresolved elements and results from preliminary analyses. Using this prompt, the agent pilot issues a request to the LLM (Step 2). The LLM responds by specifying a tool to invoke from the customized toolkit (Step 3). The agent pilot then parses the response and executes the corresponding tool with the appropriate arguments (Step 4). The tool’s output is subsequently parsed and integrated into the prompt (Step 5) for the next iteration. This loop continues iteratively until either the completer tool is executed, indicating task completion, or a predefined computational budget is reached.
@@ III-B Dynamic Prompting
WIRL adopts a reasoning-and-acting framework (i.e., ReAct [45]), iterating until either the goal is achieved or the computational budget is exhausted. In each iteration, the agent pilot parses the output of the invoked tool and incorporates the results into the prompt, which then serves as the input for the subsequent iteration. The static and dynamic sections of the prompt are described in detail below:
@@ III-B1 Role (Static)
The static section defines the agent’s role as a Senior Java Development Engineer and clearly articulates the task objective: to perform variable-level code integration based on the provided code context. Notably, this section explicitly instructs the agent to make decisions autonomously, without relying on user intervention at any point during the process.
@@ III-B2 Goals (Static)
We define three primary goals for the agent to accomplish:
@@ III-B3 Guidelines (Static)
We define a set of guidelines to which the agent should adhere.
@@ III-B4 State Description (Dynamic)
To guide the agent in utilizing the customized toolkit in a meaningful and efficient manner, we define a state machine comprising three states: initial state, insufficient context state, and sufficient context state. This mechanism is introduced because, in our preliminary experiments, we observed that the agent often engaged in aimless exploration when operating without explicit guidance.
Each state is associated with a set of available tools, as described in Section III-C, and the state description section of the prompt informs the agent of its current status. Specifically, the agent begins in the initial state, where it identifies unresolved elements and performs a preliminary analysis of the context. Based on this analysis and subsequent tool executions, the agent transitions into the insufficient context state. In this state, it iteratively invokes relevant tools to gather additional contextual information until either the context is deemed sufficient or the predefined resource budget is exhausted. Once the agent reaches the sufficient context state, it invokes the completion module (also implemented as an LLM invocation, detailed in Section III-C) to make the final recommendation.
It is important to note that while the state machine offers structured guidance, it does not enforce a fixed sequence of tool invocations. The selection and ordering of tools are left entirely to the agent’s discretion.
@@ III-B7 Output Format (Static)
To enable the agent pilot to effectively parse the output, we require that all responses from the LLM be structured in JSON format. Each JSON object must contain three mandatory fields: “thought”, “action”, and “action_input”. The “thought” field captures the agent’s reasoning and decision-making process based on the current context. The “action” field specifies the name of the tool to be invoked next, while the “action_input” field provides the necessary arguments for that tool, also formatted as a JSON object.
@@ III-C Customized Toolkit
TABLE I: Available Tools for the Agent
@@
One of the key novelties of WIRL lies in its ability to autonomously decide which tools to invoke based on the current state. The toolkit provided for WIRL (as shown in Table I) is carefully customized to facilitate the code wiring task.
@@ III-D3 Executing Tools
With the parsed tool names and corresponding arguments, the agent pilot invokes the designated tools and returns their outputs. Notably, all tool executions are performed within an isolated environment to ensure they do not interfere with the host system or the internal functioning of WIRL.
@@ IV-A Research Questions
  • RQ1: How well does WIRL perform compared with baselines?
  • RQ2: How well does WIRL perform regarding time cost?
@@ IV-B Dataset
  • As WIRL depends on compiler feedback and AST binding analysis, it requires the full project to be resolvable. Therefore, we conducted a project-wise manual inspection of the remaining dataset. The mappings were sorted in descending order based on the frequency of references to each repository, prioritizing more frequently reused code.
@@ IV-C1 SOTA Approaches
Besides that, we selected the widely adopted industrial IDE, IntelliJ IDEA [6], as one of our baseline approaches. The rename recommendation support in IDEA is both convenient and practical, and it is also suitable to address the code wiring problem.
IV-C2 LLMs
We select the following LLMs as our baselines: Distilled tiny models, including GPT-4o-mini [2], Qwen2.5-Coder-14B [3], and Qwen2.5-Coder-32B [4]. Full-parameter models, including DeepSeek-V3 [1] and Qwen-max [5].
@@ IV-D Metrics
  • #Recommendation (Rec): the number of cases where the evaluated approaches make a recommendation for the developers. Note that the number of recommendations (#recommendation) may not always match the total number of cases (#total cases).
  • #Exact Match (EM): the number of cases where the recommended code elements are identical to the ground truth (i.e., the existing names in context).
  • $EM_{Precision}$: the number of exact matches divided by the number of recommendations, i.e.,
@@
  • $EM_{Recall}$: the number of exact matches divided by the number of cases involved in the evaluation.
@@ IV-E Experimental Setup
IV-E1 Input and Output Format
@@
  • For the LLM baselines, the entire class containing the isolated code is provided as input. The isolated code, representing the segment requiring adaptation, is explicitly marked using a pair of control tokens, $<start>$ and $<end>$, to delineate the adaptation region. The complete prompt format used for the LLM baselines is available in our public repository [8].
@@
For the sake of clarity, we constrained the LLM baselines to output only the unresolved elements along with their corresponding context elements. To promote self-consistency and mitigate output variability, we set the temperature to 0. Furthermore, each evaluated approach was executed five times, and the most frequently generated output was selected as the final result, following the protocol outlined by Chen et al. [18].
As both IDEA and ExampleStack produce ranked lists of candidate substitutions, we only considered their top-ranked recommendation as the final answer to ensure a fair and consistent comparison across all evaluated approaches.
IV-E2 Implementations
To implement WIRL, we built a plugin for IntelliJ IDEA using the IntelliJ PlatForm Plugin SDK [7] and the Langchain4J [13] framework to interface with LLM APIs. While WIRL is designed to be compatible with any advanced LLMs, in this study we use Qwen2.5-Coder-14B as the backend model for two key reasons. First, if WIRL demonstrates superior performance over state-of-the-art full-parameter LLMs while relying on a smaller model, it further highlights the strength of our design. Second, models with fewer parameters typically respond faster [43, 42], making them more suitable for real-time usage in development environments.
To support code analysis and contextual understanding, we leverage the Program Structure Interface (PSI) [11], a powerful abstraction for syntactic and semantic analysis, including data flow inspection and code structure extraction.

### v2

# Wired for Reuse: Automating Context-Aware Code Adaptation in IDEs via LLM-Based Agent

@@ Abstract
Copy-paste-modify is a widespread and pragmatic practice in software development, where developers adapt reused code snippets, sourced from platforms such as Stack Overflow, GitHub, or LLM outputs, into their local codebase. A critical yet underexplored aspect of this adaptation is code wiring : the context-aware process of substituting unresolved variables in pasted code with suitable variables or expressions from the surrounding context. Existing solutions either rely on heuristic rules or historical templates, often failing to effectively utilize contextual information, despite studies showing that over half of adaptation cases are context-dependent. In this paper, we introduce WIRL, an LLM-based agent for code wiring framed as a Retrieval-Augmented Generation (RAG) infilling task. WIRL combines an LLM, a customized toolkit, and an orchestration module to identify unresolved variables, retrieve context, and perform context-aware substitutions. To balance efficiency and autonomy, the agent adopts a mixed strategy: deterministic rule-based steps for common patterns, and a state-machine-guided decision process for intelligent exploration. We evaluate WIRL on a carefully curated, high-quality dataset consisting of real-world code adaptation scenarios. Our approach achieves an exact match precision of 91.7% and a recall of 90.0%, outperforming advanced LLMs by 22.6 and 13.7 percentage points in precision and recall, respectively, and surpassing IntelliJ IDEA by 54.3 and 49.9 percentage points. These results underscore its practical utility, particularly in contexts with complex variable dependencies or multiple unresolved variables. We believe WIRL paves the way for more intelligent and context-aware developer assistance in modern IDEs.
@@ I Introduction
Copy-paste-modify is a common and inevitable practice during development. The developers frequently reuse code snippets from online programming Q&A communities, e.g., Stack Overflow [1], open-source repositories, e.g., GitHub [2], or code generation of LLMs. However, in most cases the reused code snippets need additional adaptations to integrate to the local codebase. As reported, more than 85% code snippets from Stack Overflow need to be adapted (modified) before integration into the local code [3, 4]. Variable identifiers are usually under adaptation since undeclared or conflict identifiers would lead to compilation errors or potential bugs. Consequently, code wiring, replacing unresolved variables with existing ones from the local context, is one of the most prevalent forms of adaptation. An example of code wiring practice is presented in Fig. 1. Automatic code wiring can help developers get rid of the repetitive and error-prone processes during the adaptation and concentrate on the complex business logic. However, existing approaches are designed based on either simple heuristic rules [5] or templates extracted from historical modifications [6], leaving the context not effectively leveraged although 56.1% of the adaptation cases are dependent on the surrounding context as reported by Zhang et al. [3]. Advanced code editing approaches [7, 8] are dependent on a rich history of prior edits, information that is inherently scarce in the context of code wiring. Full-parameter LLMs have demonstrated promising performance; however, their high latency renders them impractical for real-world applications. On the other hand, distilled LLMs with smaller parameter sizes tend to under perform in terms of accuracy and robustness.
@@
To this end, we present WIRL, an LLM-based agent for code Wiring through Infilling with RAG and LLMs. WIRL consists of three core components: an LLM, a customized toolkit, and an agent pilot. The customized toolkit offers three primary functionalities: (1) identifying and locating unresolved elements, (2) analyzing and collecting contextual information, and (3) infilling and recommending suitable substitutions. The agent pilot is responsible not only for initializing the prompt but also for coordinating the interaction between the LLM agent and the toolkit. It parses the LLM’s output, updates the prompt dynamically, and invokes the appropriate tools as needed. With a dynamically updated prompt, WIRL incrementally gathers relevant context by invoking appropriate tools until the agent determines that sufficient information has been collected to make a final recommendation. To fully leverage the capabilities of LLMs, we reformulate the adaptation task as a retrieval-augmented generation (RAG)-based infilling task for the unresolved elements, framing it as a code completion problem which aligns more naturally with the strengths of LLMs. Empirical evidence from Zhang et al. [9] supports this reformulation, showing that LLM performance on code snippet adaptation tasks lags behind their performance on code completion and generation tasks. Moreover, by equipping with the RAG-based infilling strategy, even distilled LLMs with smaller parameter sizes can outperform full-parameter LLMs, thereby meeting both the latency and accuracy requirements of real-world software development scenarios. To enhance the efficiency of WIRL, we adopt a hybrid execution mode for the agent, wherein essential steps are extracted and executed initially without invoking the LLM. In addition to this optimization, we introduce a state machine to guide the agent’s exploration process, thereby reducing unnecessary or ineffective attempts and improving overall execution efficiency.
@@
  • We develop a customized toolkit that provides essential functionalities for precise and efficient code adaptation.
@@
The rest of this paper is structured as follows. Section II motivates this study. Section III illustrates the design of WIRL. Section IV presents experimental details and results. Section V first explores the broader implications and robustness of our findings through analysis of applicability and versatility. It then discusses potential threats to validity and the inherent limitations of our study. Section VI discusses the related works, and Section VII concludes this paper.
@@ II Motivating Example
We first present the problem formulation, which serves as an important foundation of our approach. Assume the set of unresolved elements is denoted by $\mathcal{U}={u_{i}}$ and the set of context elements by $\mathcal{C}={c_{i}}$. The code wiring task can be formally defined as identifying an injective mapping $\mathcal{M}:\mathcal{U}\rightarrow\mathcal{C}$, where each $u_{i}\in\mathcal{U}$ is mapped to a distinct $c_{i}\in\mathcal{C}$. In this work, however, we reformulate code wiring as a retrieval-augmented generation (RAG)-based infilling task. Specifically, each $u_{i}\in\mathcal{U}$ is replaced with a placeholder in the isolated code. The objective then becomes to infill these placeholders using the retrieved context information, including the variable names themselves, which is motivated by the observation that literal similarity often provides valuable contextual cues. This reformulation is grounded in the insight that code completion is more naturally aligned with the capabilities of LLMs than explicit code adaptation. This design choice is further supported by recent findings by Zhang et al. [9], which highlight the superior performance of LLMs in code completion tasks relative to direct adaptation scenarios.
@@
Figure 2: Overview of WIRL
@@ III-A Overview
The overview of WIRL is illustrated in Fig. 2. WIRL consists of three core components: an LLM, a customized toolkit containing three types of tools, i.e., locator, collector, and completer, and an agent pilot, which orchestrates communication between the LLM and the toolkit by parsing outputs, updating prompts, and executing tool invocations. Given a piece of isolated code, the agent pilot begins by initializing the prompt (Step 1). This prompt includes system roles, task descriptions, and instructions on how the task should be addressed using the available tools. It also embeds the location of unresolved elements and results from preliminary analyses. Using this prompt, the agent pilot issues a request to the LLM (Step 2). The LLM responds by specifying a tool to invoke from the customized toolkit (Step 3). The agent pilot then parses the response and executes the corresponding tool with the appropriate arguments (Step 4). The tool’s output is subsequently parsed and integrated into the prompt (Step 5) for the next iteration. This loop continues iteratively until either the completer tool is executed, or a predefined computational budget is reached. The customized toolkit, implemented by the proposed approach, depends on high-level abstractions like compilers, AST parsers, and LLMs.
@@ III-B Dynamic Prompting
Adopting the ReAct framework [13], WIRL iteratively reasons and acts until the goal is met or the budget is exceeded. In each cycle, the agent pilot parses tool outputs and integrates them into the prompt for the next iteration, which consists of both static and dynamic sections.
@@ III-B1 Role (Static)
The static section defines the agent’s role as a Senior Java Development Engineer, tasked with autonomous, variable-level code integration based on the provided context, explicitly prohibiting any user intervention.
@@ III-B2 Goals (Static)
Three primary goals to accomplish:
@@ III-B3 Guidelines (Static)
Guidelines the agent should adhere:
@@ III-B4 State Description (Dynamic)
To guide the agent in utilizing the customized toolkit in a meaningful and efficient manner, we define a state machine comprising three states: initial state, insufficient context state, and sufficient context state. Each state is associated with a set of available tools, as described in Section III-C, and the state description section of the prompt informs the agent of its current status. This mechanism is introduced because, in our preliminary experiments, we observed that the agent often engaged in aimless exploration when operating without explicit guidance. The transition rules between each state are presented in Table I. The agent begins in the initial state, where it identifies unresolved elements and performs a preliminary analysis of the context. It then transitions to the insufficient context state to iteratively gather necessary contextual information. The agent moves to the sufficient context state once enough information has been gathered or a predefined resource budget (i.e., iteration limit) is exhausted. In this state, it invokes the completer to generate the recommendation before terminating (done state). This structured workflow, with its built-in budget to handle errors and prevent infinite loops, ensures a focused and efficient decision-making process. It is important to note that while the state machine offers structured guidance, it does not enforce a fixed sequence of tool invocations. The selection and ordering of tools are left entirely to the agent’s discretion.
TABLE I: Transition Rules of State Machine
From State Transition Rules To State
Initial Initialization Finished Insufficient Context
Insufficient Context
Budget Exhausted or
Context Deemed Sufficient
Sufficient Context
Sufficient Context Completer Executed Done
@@ III-B7 Output Format (Static)
To enable the agent pilot to effectively parse the output, we require all responses from the LLM be structured in JSON format. Each JSON object must contain three mandatory fields: “thought”, “action”, and “action_input”. The “thought” field captures the agent’s reasoning and decision-making process based on the current context. The “action” field specifies the tools’ name to be invoked next, while the “action_input” field provides necessary arguments for that tool, also formatted as a JSON object.
@@ III-C Customized Toolkit
TABLE II: Available Tools for the Agent
@@
One of the key novelties of WIRL is its ability to autonomously decide which tools to invoke based on the current state. The toolkit provided for WIRL (as shown in Table II) is carefully customized to facilitate the code wiring task.
@@ III-D3 Executing Tools
With the parsed tool names and corresponding arguments, the agent pilot invokes the designated tools and returns their outputs. Notably, all tool executions are performed within an isolated environment to ensure they do not interfere with the internal functioning of WIRL.
@@ IV-A Research Questions
  • RQ1: How well does WIRL perform compared against baselines?
  • RQ2: How well does WIRL perform regarding latency?
@@
RQ4 quantifies the contribution of each of WIRL’s design components, namely the Locator, Collector, and Completer tools, as well as its architectural design, to the system’s overall performance. Understanding the individual impact of these elements is crucial for validating the novelty of our approach and justifying the design of the WIRL framework.
@@ IV-B Dataset
  • As WIRL depends on compiler feedback and AST binding analysis, it requires the project to be resolvable. Therefore, we conducted a project-wise manual inspection of the remaining dataset. The mappings were sorted in descending order based on the frequency of references to each repository, prioritizing more frequently reused code.
@@ IV-C1 SOTA Approaches
Since code editing task is potential to solve the code wiring problem, we include two state-of-the-art approaches in code editing, e.g., GrACE [7] and CoEdPilot [8].
@@
Besides that, we selected the widely adopted industrial IDE, IntelliJ IDEA [5], as one of our baselines. The renaming recommendation support in IDEA is both practical and suitable to address the code wiring problem.
IV-C2 Raw LLMs
We take the following raw LLMs as our baselines: Distilled models, i.e., GPT-4o-mini [17], Qwen2.5-Coder-14B [18], and Qwen2.5-Coder-32B [19]; Full-parameter models, i.e., DeepSeek-V3 [20] and Qwen-max [21].
@@ IV-D Metrics
  • #Recommendation (#Rec): the number of cases where the evaluated approaches make a recommendation for the developers. Note that the number of recommendations (#recommendation) may not always match the total number of cases (#total cases).
  • #Exact Match (#EM): the number of cases where the recommended code elements are identical to the ground truth (i.e., the existing names in context).
  • $EM_{Precision}$ ($EM_{P}$): the number of exact matches divided by the number of recommendations, i.e.,
@@
  • $EM_{Recall}$ ($EM_{R}$): the number of exact matches divided by the number of cases involved in the evaluation.
@@ IV-E Experimental Setup
IV-E1 Input and Output Format for LLMs
@@
  • For the LLM baselines, the entire class containing the isolated code is provided as input. The isolated code, representing the segment requiring adaptation, is explicitly marked using a pair of control tokens, <start><start> and <end><end>, to delineate the adaptation region. The processed data format and the complete prompt format used for the LLM baselines are available in our public repository [23].
@@
For the sake of clarity, we constrained the LLM baselines to output only the unresolved elements along with their corresponding context elements.
IV-E2 LLM Configurations
To promote self-consistency and mitigate output variability, we set the temperature to 0. The top-p is set to 0.2, and the other parameters are set to default. Furthermore, each evaluated approach was executed five times, and the most frequently generated output was selected as the final result, following the protocol outlined by Chen et al. [24].
IV-E3 Baseline Configurations
As IDEA, ExampleStack, and CoEdPilot produce ranked lists of candidate substitutions, we only considered their top-ranked recommendation as the final answer to ensure a fair and consistent comparison across all evaluated approaches. For CoEdPilot, we also equipped it with our Locator module to maximize its performance.
Due to the deprecation of the code-davinci-002 model, which was originally used in the zero-shot implementation of GrACE, we sought a suitable replacement. OpenAI officially discontinued support for the Codex API, including code-davinci-002, on March 23, 2023, and recommended transitioning to more advanced models. To this end, we opted to use a more advanced model, GPT-4o, for our implementation of GrACE.
IV-E4 Implementations
To implement WIRL, we built a plugin for IntelliJ IDEA using the IntelliJ PlatForm Plugin SDK [25] and the Langchain4J [26] framework to interface with LLM APIs. While WIRL is designed to be compatible with any advanced LLMs, in this study we use Qwen2.5-Coder-14B as the backend model for two key reasons. First, if WIRL demonstrates superior performance over state-of-the-art full-parameter LLMs while relying on a smaller model, it further highlights the strength of our design. Second, models with fewer parameters typically respond faster [27, 28], making them more suitable for real-time usage in development environments. In our current implementation, we utilize the Program Structure Interface (PSI) [29], a powerful toolkit for syntactic and semantic analysis, to gather compiler information for the Locator and perform AST analysis for the Collector.